OpenStudy (anonymous):

write the slope-intercept form of the equation of the line which passes through the points (1,6) and (2,9). a. y=2x+3 b.y=3x+3 c.y=6x+9 d.y=3x+6

OpenStudy (anonymous):

Slope is (9-6)/(2-1). So your slope is 3. Plug a point into point slope form and you will get y-6=3(x-1). Distribute: y-6=3x-3. Move the 6 over: y=3x+3